40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8676     Entry time: Wed Jun 5 10:46:43 2013
Author: Gautam 
Type: Update 
Category: General 
Subject: Cable re-routing at 1Y4 

 There were 4 cables running over the front side of rack 1Y4 such that the front door could not be closed. I re-routed them (one at a time) through the opening on the top of the rack. The concerned channels were

  • Green refl mon
  • Err mon
  • Pzt out (temp) (has been marked "Door Damaged BNC")
  • Laser temp ctrl

Before and after pics attached.

Before

.before.JPG

After

after.JPG

 

 

 

ELOG V3.1.3-